# Distilling Event Sequence Knowledge From Large Language Models

### v1

@@ 4.3 Human Evaluation of Cause-Effect Prediction Accuracy of LLMs
In addition to this automated evaluation of our LLM-generated sequences, we carry out additional evaluation of our method on an existing popular knowledge graph – ATOMIC Hwang et al. (2021) and conduct human evaluation on a sample of data to support our approach. ATOMIC consists of event-centered pairs of instances of the form IsAfter (Y comes after X) and Causes (X causes Y). We leverage these instances as input prompts to the model and use human annotators to evaluate outputs. Specifically, we show human annotators anonymized model outputs and true references and elicit their preferences given a trigger event. We find that in an overwhelming majority of the cases, the models generate output that is semantically equivalent to the reference (even though there is no direct lexical alignment), or output that the humans prefer over the true reference. We provide example outputs of various evaluator models in the supplementary material (appendix).
4.4 Evaluation of Event Models
Apart from a qualitative manual analysis of the mined patterns, we follow Bhattacharjya et al. (2022) and use a metric that evaluates the quality of the learned model with respect to the data. Specifically, we use negative log loss as the evaluation metric, which is a popular metric for probabilistic prediction Bishop (2006). We posit that higher-quality event sequences result in better prediction ability of the models, and so better negative log loss scores for the models. Our results show that this is a valid assumption, as larger models that are expected to produce higher-quality sequences yield models with better prediction ability.
($\downarrow$) Data Generator Model BSuMMs OSuMMs LSTM
Flan-T5-Large (783M) Chung et al. (2022) -63.49 -84.29 -109.28
Flan-T5-XL (3B) -63.20 -92.65 -121.23
Open-Research LLaMA (3B) Touvron et al. (2023) -110.57 -101.29 -190.68
Flan-T5-XXL (11B) -57.99 -78.64 -108.89
5 Experimental Investigation and Results
We conducted a three-pronged quantitative and qualitative evaluation of our results. First, we evaluated LLM-generated event sequences in absolute terms through standard traditional metrics – precision, recall, and F1 score. Second, we gauge how probabilistic models like SuMMs fare when trained on LLM-generated data. Finally, we look at some qualitative examples of newly identified high support sequences (Appendix LABEL:sec_apx:examples) not already present in our base knowledge graph i.e. Wikidata.
5.1 Generated Event Sequences
We generate event sequences via LLMs through iterative in-context prompting (Figure 3). Table 1 summarizes our results from these experiments. We find that Flan-T5-XXL consistently outperforms its smaller counterparts, indicating that larger models generally fare better and benefit from their larger size (i.e. number of parameters). We posit this may also indicate that using larger models may continue to yield performance gains.
@@ 5.1 Generated Event Sequences
One of the responses above is the model (Flan-T5-XXL) output, while the other is the true reference. The human evaluators are then asked to answer the following questions–
@@
Based on the responses from human evaluators [^5] We observe strong inter-rater agreement with a Fleiss kappa, $\kappa=0.81$; conflicting response labels were aggregated through majority vote., we observe that humans found $65.82\%$ of response pairs functionally equivalent. That is, even though not lexically aligned, they meant the same thing. In $27.64\%$ of the instances the humans preferred the model-generated event instance over the true reference. In only $6.55\%$ of instances did the humans prefer the true reference over the model-generated output. [^6] Additional details on these experiments are available in the supplementary material (appendix). These results reinforce our underlying assertion that LLMs are capable of event-centered commonsense reasoning.
Precision Evaluator Models
@@
We leverage our models for dual purposes – for generating and evaluating event sequences. Owing to this inherent conflict, we find it prudent to independently assess these evaluator models. To ensure robustness of our results, we use multiple evaluator models (Table 1) and find no significant difference between evaluated precision across different models used as evaluators.
5.2 Probabilistic Prediction via SuMMs and LSTM
Following the evaluation strategy implemented in Bhattacharjya et al. (2022), we focus on individual labels of interest and conduct an evaluation around probabilistic prediction. Consequently, we select negative log loss as the evaluation metric. Table 2 summarizes our results on the test set for both BSuMMs and OSuMMs, along with a simple LSTM baseline. We observe that models trained on event sequences generated from larger models (e.g. Flan-T5-XXL) fare better than the ones generated from their smaller counterparts (e.g. Flan-T5-Large). We treat this observation as a proxy for generated event sequence quality. That is, better quality sequences lead to better predictive models.
5.3 Qualitative Analysis
Figure 1 shows examples of mined patterns along with their support levels, as well as the probabilistic association between a number of influencing events from the output of BSuMMs. Appendix LABEL:sec_apx:examples provides two examples of newly mined sequences derived from pattern mining over LLM-generated sequences, and an additional two examples of influencing sets derived from the application of SuMMs. For each example, we show evidence supporting the accuracy of the extracted knowledge. We provide additional qualitative examples in supplementary material. Overall, these results show the high quality and usefulness of the distilled event sequence knowledge.
5.4 Error Analysis
We now describe two of the most common recurring errors in the LLM-generated outputs and SuMMs’ discovered influence sets.
No Sequences Generated
@@ 5.4 Error Analysis
Out of the $202$ unique input event-triggers in our data, $8$ event-triggers yielded sequences of length $1$, i.e., no sequences. This means that the underlying model could not select a relevant event concept from the provided vocabulary that might follow the given trigger.
Illogical Influencing Sets
@@
Due to the choice of the vocabulary of limited event concepts, we often find that the identified influencing sets (i.e., parent events) for some event concepts may not be correct or, at times, completely illogical. Qualitatively, we observe that this occurs with rare event types that do not appear often either in the underlying knowledge graph, or in the LLM-generated set of event sequences. Consequently, SuMMs fail to identify logical influencing sets. For example, for the event type United States Presidential Impeachment, BSuMM model identified unequal treaty as the influencing event.
6 Related Work
News Event Analysis
The primary applications of our work are around news event analysis and forecasting.  Zhao (2021) presents a taxonomy of different flavors of event prediction in the literature. Our target event prediction applications fall under the “Semantic Prediction” category, with time and location details not being of interest, and the primary goal being the prediction of “event profiles" such as event types. Seminal work in this area is the work of Radinsky et al. (2012) where causal relations between past events are extracted from text and then generalized to make predictions. Some recent work has explored the use of graph sequence mining Cekinel and Karagoz (2022) as a mechanism of extracting useful relations from large and noisy outputs of extraction. The application of building event sequence models over such outputs is largely unexplored. Our work explores an alternative approach of event sequence generation that is capable of generating longer, and potentially higher-quality sequences.
Event Sequence Extraction from Text
There is a wealth of literature on different methods of extracting sequences from textual corpora. Norambuena et al. (2023) present a comprehensive survey of automated methods of narrative extraction. Narratives contain several elements including events, participants (actors/protagonists), time, and space. The task of event detection itself is a highly challenging task and the topic of extensive research Chen et al. (2021); Xiang and Wang (2019), which has its root in the Topic Detection and Tracking (TDT) task Santana et al. (2023). This line of work started out as a DARPA-sponsored initiative with the same name Allan et al. (1998). Another closely related task is news timeline summarization Gholipour Ghalandari and Ifrim (2020). While pattern mining algorithms have been applied to the output of such extractions, e.g. for creation of “domain templates" Filatova et al. (2006), we are not aware of any attempts to use the extractions to construct event models for analysis or prediction.
Sequential Pattern Mining and Event Sequence Models
Sequential pattern mining and related approaches have been the subject of extensive research Mannila et al. (1997); Mabroukeh and Ezeife (2010); Mooney and Roddick (2013); Fournier-Viger et al. (2017). These algorithms take in a set of sequences (or “sequential records") with a set of unique events (or “items") and often a “minimum support" threshold, and return as output a ranked set of all frequent sequences in a given sequence collection (or database) meeting the minimum support threshold. There is also a long line of work on statistical and probabilistic models for analysis of different kinds of sequences. Our focus in this paper is on multivariate event sequences, i.e., sequences of various event types without timestamps. Markov models for sequences Raftery (1985); Begleiter et al. (2004) and long short-term memory (LSTM) Hochreiter and Schmidhuber (1997) models are examples of prediction models. We leverage a more recent family of models – Summary Markov models Bhattacharjya et al. (2022) – for some of the experiments in this paper that involve analyzing generated event sequences.
7 Conclusions
@@
We evaluated the capabilities of large language models for the task of producing event sequences from concepts derived from a knowledge graph. We found that, through careful elicitation, LLMs are capable of producing high-quality event sequences that can be used for the downstream tasks of pattern mining and the construction of probabilistic event models such as SuMMs. We also proposed methods for effectively evaluating the outputs of LLMs with sparse reference sets and deriving useful signals for model selection through the use of existing traditional metrics. Our hope is that this work demonstrates the utility of large language models for extracting event-related information and helps researchers better navigate the complex interaction between event-related entities. While we have demonstrated some use-cases of the resulting dataset of event sequences (mining logical rules and patterns and extracting influencing event types), resulting datasets in this work and in other domains may themselves be a useful resource for researchers as one may draw upon them to inform or justify decision-making.

### v3

@@ 4.1 Human Evaluation of Cause-Effect Prediction Accuracy of LLMs
To quantify how well model outputs correlate with human assessments, we first conduct a small-scale human evaluation on a different, independent event-commonsense reasoning dataset: ATOMIC [21]. ATOMIC consists of event-centered pairs of instances of the form IsAfter (Y comes after X) and Causes (X causes Y). We use these instances as input prompts to the model and then ask human annotators to evaluate model outputs. Specifically, we show human annotators anonymized model outputs and true references and elicit their preferences given a trigger event.
@@
One of the responses above is the LLM output, while the other is the true reference. The human evaluators are then asked to answer the following questions–
@@
We find that in an overwhelming majority of the cases, the models generate output that is semantically equivalent to the reference (even though there is no direct lexical alignment), or output that the humans prefer over the true reference. Based on the responses from human evaluators [^2] We observe strong inter-rater agreement with a Fleiss kappa, $\kappa=0.81$; conflicting response labels were aggregated through majority vote., we observe that humans found $65.82\%$ of response pairs functionally equivalent. That is, even though not lexically aligned, they meant the same thing. In $27.64\%$ of the instances the humans preferred the model-generated event instance over the true reference. In only $6.55\%$ of instances did the humans prefer the true reference over the model-generated output. [^3] Additional details on these experiments are available in the supplementary material (Appendix). These results reinforce our underlying assertion that LLMs are capable of event-centered reasoning, and therefore could produce high-quality event sequence collections.
4.2 Evaluation of Recall
Despite the sparsity of causal relations in Wikidata, one can still reasonably estimate recall through pairwise comparisons of generated event classes to the existing causal relations in Wikidata. We build a reference set of causal relations in Wikidata by curating a list of all the pairs of event concepts that have any of the several causal relations444www.wikidata.org/wiki/Wikidata:List_of_properties/causality in Wikidata in any direction, including has_cause (P828), has_immediate_cause (P1478), has_contributing_factor (P1479), and has_effect(P1542). Our results show that this is an effective mechanism of comparing recall across methods, as larger models that are expected to have better accuracy yield higher recall scores.
4.3 Evaluation of Precision
To overcome sparsity in reference data, prior work has generally relied on human evaluations [10] for estimating precision, which entails looking at pairs of events and asking annotators whether or not the events in question have a causal relationship. Such a process for potentially thousands of event pairs (like in our case) can be very cost prohibitive. Furthermore, recent research [46, 19, 47] indicates that pre-trained language models themselves might outperform lay human annotators such as those found on Amazon Mechanical Turk. For instance, [19] demonstrated that labeling data under a few-shot chain-of-thought prompt (“explain-then-annotate” setting) surpasses crowdworker annotations on relevance assessments. Following these results, we use LLMs for evaluating precision and for selecting the most optimal model. We propose evaluating precision for model selection as a binary classification task. Given two events $e_{1},e_{2}$, an evaluator model must evaluate whether $e_{1}$ reasonably leads to $e_{2}$ under a has_cause or has_effect relationship. To do this, we start with a large instruction-tuned model (in our case, Flan-T5-XXL ($11$B) [11]) and adopt instructional in-context few shot prompting to classify whether or not $e_{1}$ reasonably leads to $e_{2}$, and for the model to provide a justification for its results. While we hypothesize that such an approach may yield noisy results, a small manual assessment as well as our results comparing different models with different evaluator models indicate that this approach yields reliable results for comparing different models, and a reasonable estimate of the overall precision of the model. Note that the evaluator performs only a simple binary classification task that has a very high accuracy even in smaller models.
4.4 Settings
We performed all our model inference related experiments on two NVIDIA V100 GPUs. We used the Huggingface library v4.26.14.26.14.26.14.26.1 [40] and publicly available model checkpoints.555huggingface.co/docs/transformers/model_doc/flan-t5 Classical pattern mining algorithms (GSP, SPADE) were implemented in Python through the use of spmf666pypi.org/project/spmf/ library v1.41.41.41.4. For event sequence generation, we use in-context learning under zero shot settings for all LLMs, with top-k sampling ($k=50$, in conjunction with top-p, where $p=0.95$). To identify influencing events from Summary Markov Models (SuMMs), we use implementations from [6] and split the dataset (generated from LLMs) into train/dev/test sets ($70\%/15\%/15\%$) to generate results reported in Table 2. BSuMMs and OSuMMs were learned with hyperparameters of $\alpha=0.1$, $\gamma=0.5$ and a look-back ($\kappa$) window of $4$.
4.5 Results
Table 1 summarizes our results from these experiments. While prior work as proven the concept of LLM-as-a-judge [47] when much larger LLMs like GPT-4 are used, here we use less resource-intensive LLMs not for evaluation of the absolute quality of the generated sequences, but for comparison of the relative performance of models
@@ 4.5 Results
Since we use our models for dual purposes – for generating and evaluating event sequences, owing to this inherent conflict we find it prudent to independently assess these evaluator models. To ensure robustness of our results, we use multiple evaluator models and find no significant difference between evaluated precision across different models used as evaluators with the largest model performing marginally better as a precision evaluator.
Precision Evaluator Model (P) R F-1
Flan-T5-Large Flan-T5-XL Flan-T5-XXL
Flan-T5-Large ($783$M) [11] 0.73 0.72 0.72 0.47 0.57
Flan-T5-XL (3B) 0.75 0.75 0.78 0.49 0.60
Open-Research LLaMA ($3$B) [35] 0.70 0.69 0.72 0.45 0.55
Flan-T5-XXL ($11$B) 0.79 0.79 0.81 0.54 0.65
Table 1: Evaluation of LLM-generated sequences. For the purpose of evaluating recall (R), we count event pairs (e1, e2) if such a pair exists in Wikidata. For evaluating precision (P), we treat correctness of all generated event pairs (e1, e2) as a standard classification task. Best-performing model scores are in bold.
4.6 Discussion
@@
We observed in our analysis that some event triggers did not produce any sequences. Out of the $202$ unique input event triggers in our data, $8$ event triggers yielded sequences of length $1$, i.e., no sequences. This means that the underlying model could not select a relevant event concept from the provided vocabulary that might follow the given trigger.
5 Knowledge Distillation Through Event Sequence Analysis
Given a high-quality collection of event sequences generated by utilizing LLMs, we use pattern mining to discover high-support sequence patterns not directly derivable from the knowledge graph, and learn probabilistic models to discover complex event sequence rules.
5.1 Mining New Patterns
In order to derive new, unseen relationships (has_cause or has_effect) between the extracted event classes, we use the generated event sequence collection followed by classical frequent itemset mining algorithms like GSP [34] and SPADE [43] to derive new high support patterns. This aspect of the overall workflow is illustrated in the middle section of Figure 1, where high-support patterns are mined from event sequences.
The classical pattern mining algorithms we use to mine for new event patterns are both Apriori-based approaches. Under both methods, given two sequences of the same event concept class
$\alpha=<a_{1},a_{2},..,a_{n}>\text{ and }\beta=<b_{1},b_{2},..,b_{n}>$ (2)
$\alpha$ is a subsequence of $\beta$ denoted as $\alpha\subseteq\beta$ iff there exists a set of values $1\leq j_{1}<j_{2}<...<j_{n}\leq m$ such that $a_{1}\subseteq b_{j1},a_{2}\subseteq b_{j2},...,a_{n}\subseteq b_{jn},$; then $\beta$ is a supersequence of $\alpha$. Given multiple such sequences and a support threshold, the task here is to find a set of frequent event subsequences. GSP is depicted in Algorithm 1). SPADE uses a "vertical database format", which stores sequences as lists of itemsets associated with their IDs (referred to as “id-lists"), which allows faster support counting. The algorithm organizes the search space into equivalence classes, and avoids multiple scans of the sequences.
Algorithm 1 Generalized Sequential Pattern (GSP) Mining Algorithm
1:Input: Database of sequences $D$, minimum support threshold min_sup
2:Output: Set of all sequential patterns $SP$
3:$SP\leftarrow\emptyset$
4:$C_{1}\leftarrow$ set of all individual items in $D$ that meet min_sup
5:$L_{1}\leftarrow$ filter candidates in $C_{1}$ by min_sup
6:$k\leftarrow 2$
7:while $L_{k-1}\neq\emptyset$ do
8:     $C_{k}\leftarrow$ generate_candidates($L_{k-1}$)
9:     for each sequence $s\in D$ do
10:         for each candidate $c\in C_{k}$ do
11:              if $c$ is contained in $s$ then
12:                  increment support count of $c$
13:              end if
14:         end for
15:     end for
16:     $L_{k}\leftarrow$ filter candidates in $C_{k}$ by min_sup
17:     $SP\leftarrow SP\cup L_{k}$
18:     $k\leftarrow k+1$
19:end while
20:return $SP$
5.1.1 Examples of Mined Patterns
Following are example outputs from application of GSP on LLM outputs.
Mined Pattern Example 1
Pattern: Civil Disorder (Q686984) $\rightarrow$ Democratization (Q1064441) $\rightarrow$ Energy Crises (Q8413663) Support: 5
The pattern above occurs with a support value of $5$ i.e. at least supported by $5$ super-sequences. Empirically, we can find support for such a pattern in history.777wikipedia.org/wiki/South_African_energy_crisis In 1994, the country of South Africa democratized post civil disorder. This led to an increased energy demand over the following decade, eventually culminating in a full blown energy crises starting 2003.
Mined Pattern Example 2
Pattern: Famine (Q168247) $\rightarrow$ Refugee Crises (Q20898283) $\rightarrow$ Post Traumatic Stress Disorder(Q202387)
Support: 5
The pattern above again occurs with a support value of $5$ i.e. at least supported by $5$ super-sequences. Similar to the previous example, evidence supporting such an event tranisiton can be found in real life.888wikipedia.org/wiki/Great_Famine_(Ireland) During the Great Irish Famine people were forced to relocate and flee Ireland causing a refugee crises. A great number of these individuals suffered from mental health crises (e.g. PTSD) due to the events directly associated with the famine.
5.2 Identifying Influencing Sets through Summary Markov Models
Learning about potential influencing events that lead to a given event type in a large set of event sequences is an important aspect of analyzing multivariate event sequences, i.e. events without time stamps, like the ones generated by our models. Classic $k_{th}$ order Markov chains capture these dynamics by modeling the probability of observing a particular event type given the preceding $k$ events in-sequence. The recent family of summary Markov models (SuMMs) [6] generalize other well known Markov models for event sequences by leveraging a function that summarizes historical event occurrences, and identifying the subset of event types that affect the probability of occurrence of event types of interest; this forms the influencing set.
We use the LLM-generated sequences to learn two types of SuMMs: binary SuMMs (BSuMMs) and ordinal SuMMs (OSuMMs). In BSuMMs, it is only the presence or absence of an event in the relevant history that has an effect on the occurrence of other events, while in OSuMMs the order of the events is also taken into account. We refer the reader to Sections 3.3 and 3.4 in [6] for complete formal definitions and methods for learning SuMMs over event sequence collections. Briefly, given a subset of event labels of interest $\mathbf{X}\in\mathcal{L}$ and a set of parameters $\Theta_{X}=\{\theta_{x|h}\}$, where $\theta_{x|h}$ is the probability of a given event label $x\in\mathbf{X}$ occurring at any position in the sequence given the historical event occurrences $h$, influencing and non-influencing event sets can be formally defined as label sets U = $\mathcal{L}\setminus$ U are influencing sets for event labels $\mathbf{X}$ such that they minimally determine the probability of observing any particular label of interest $x_{i}$ for a given position $i$ in the sequence.
5.2.1 Evaluation: SuMMs vs LSTMs
Following the evaluation strategy implemented in [6], we focus on individual labels of interest and conduct an evaluation around probabilistic prediction. Consequently, we select negative log loss as the evaluation metric. Table 2 summarizes our results on the test set for both BSuMMs and OSuMMs, along with a simple LSTM baseline. We observe that models trained on event sequences generated from larger models (e.g. Flan-T5-XXL) fare better than the ones generated from their smaller counterparts (e.g. Flan-T5-Large). We treat this observation as a proxy for generated event sequence quality. That is, better quality sequences lead to better predictive models.
($\downarrow$) Data Generator Model BSuMMs OSuMMs LSTM
Flan-T5-Large (783M) [11] -63.49 -84.29 -109.28
Flan-T5-XL (3B) -63.20 -92.65 -121.23
Open-Research LLaMA (3B) [35] -110.57 -101.29 -190.68
Flan-T5-XXL (11B) -57.99 -78.64 -108.89
5.2.2 Qualitative Assessment
Figure 1 shows examples of learned influencing sets using BSuMMs for refugee crisis and mass migration events. Two events mass migration and famine are identified as influencing events for refugee crisis. The model indicates that the occurrence of both mass migration and famine together has a $0.91$ probability of resulting in refugee crisis as a part of a sequence of events. On the other hand the occurrence of mass migration in the absence of famine has a $0.31$ probability of resulting in refugee crisis. BSuMMs and OSuMMs deploy a greedy score-based forward and backward search strategy to efficiently discover the minimal influencing sets. Overall, the discovery of influencing sets from LLM-generated data provides a mechanism of distilling complex symbolic knowledge from the output of LLMs.
We provide two more examples below of influencing sets derived from the application of SuMMs. For each example, we show evidence supporting the accuracy of the extracted knowledge. Overall, these results show the high quality and usefulness of the distilled event sequence knowledge.
Influencing Set Example 1
The following example identifies “Hate Crimes” as a predecessor to the occurrence of “Civil Disorder” related events.
X: Civil Disorder (Q686984)
Parent: Hate Crime (Q459409)
  • $P(\text{{Civil Disorder}}|\text{{NO Hate Crime}})=0.12$
  • $P(\text{{Civil Disorder}}|\text{{Hate Crime}})=0.55$
In the example above, we see that the likelihood of a civil disorder is greatly influenced by the occurrence of a hate crime.
Influencing Set Example 2
The following example identifies “Disease Outbreaks” and “Lockdowns” as precursors to the institution of “Travel Restrictions”.
X:Travel Restriction (Q87745167)
Parents: Disease Outbreak (Q3241045), Lockdown (Q6665312)
[…]
Quite intuitively, we see here the likelihood of Travel Restrictions directly correlate with the institution of Disease Outbreaks and Lockdowns (since the latter have been lately associated with the former).
5.2.3 Error Analysis
@@
We observed that in some cases, the identified influencing setsfor some event concepts were not correct or, at times, illogical. Qualitatively, we observe that this occurs with rare event types that do not appear often either in the underlying knowledge graph, or in the LLM-generated set of event sequences. Consequently, SuMMs fail to identify logical influencing sets. For example, for the event type United States Presidential Impeachment, BSuMM model identified unequal treaty as the influencing event.
6 Conclusions and Future Work
In this paper, we presented methods of distilling event sequence knowledge from large language models. We first presented an approach of generating high-quality event sequences through knowledge-guided generation using LLMs. We then conducted a three-pronged quantitative and qualitative evaluation of our results. First, we evaluated LLM-generated event sequences in absolute terms through manual evaluation and automated evaluation using standard accuracy metrics. Second, we looked at some qualitative examples of newly identified high support sequences not already present in our base knowledge graph i.e. Wikidata. Finally, we gauge how probabilistic models like SuMMs fare when trained on LLM-generated data. We found that through our carefully-designed elicitation procedure, LLMs are capable of producing high-quality event sequences that can be used for the downstream tasks of pattern mining and the construction of probabilistic event models such as SuMMs.
@@ 6 Conclusions and Future Work
This work demonstrates the utility of large language models for extracting event-related information and helps researchers better navigate the complex interaction between event-related entities. While we have demonstrated some use-cases for the resulting dataset of event sequences – mining logical rules/patterns and extracting influencing event types – the resulting datasets in this work and those from other domains could themselves be a useful resource for researchers, as one may leverage them suitably to inform or justify decision-making. We make our code and datasets publicly available for future research. Here, we outline a number of directions for future work:
  • In this paper, we used Wikidata as our base source of knowledge. The event concepts found on Wikidata often appear in the pre-training datasets of large language models. We did not consider more complex, domain-specific datasets of non-timestamped events (e.g. Healthcare, Finance). Applying LLMs to generate domain-specific events may require considerable amounts of data to fine-tune these models, and collecting such supervision to train these models at scale may be prohibitively expensive.
  • A key element of our evaluation strategy (precision) involves the use of the same language models that were used to generate the sequences being evaluated in the first place. Through qualitative examples and past research on using LLMs as annotators, we observe that such a strategy yields a noisy but useful signal for evaluating model performance. However, an ideal and accurate evaluation of model outputs should involve the use of human annotators and a validation of the use of larger LLMs as a part of LLMs-as-a-judge [47] evaluation strategy. Furthermore, our use of smaller models was justified in part due to the simplicity of the precision evaluation task, which can be posed as a binary classification task. It will be interesting to evaluate the correctness of post-hoc explanations generated by models for their corresponding output during precision evaluation, as such explanations are key to the usefulness of the distilled knowledge, particularly in critical decision support applications.
  • In the context of our domain of interest in this paper (news event analysis), an exciting avenue for future research is to study the effect of utilizing structured knowledge distilled from LLMs in a pipeline for future event prediction, and a comparison with the performance of human forecasters. Recent work [17] claims that LLMs, without symbolic knowledge elicitation, come close and in some cases surpass the ability of a “crowd aggregate of competitive forecasters" in making accurate forecasts in the form of answering forecast questions on online platforms such as the Good Judgment Open ((https://www.gjopen.com)). Two research questions arise: 1) Could a neurosymbolic approach utilizing distilled symbolic knowledge complement or outperform a purely LLM-based solution? 2) Would distilled high-quality structured knowledge provide better explainability and therefore be a more reliable tool as a part of a human-AI collaboration mechanism for event forecasting?
Supplemental Material Statement:
Our supplementary material (included as a zip file in our submission) includes code and the prompts used for event sequence generation and benchmarking, as well as the base KG, and sample output files. README.md has all the details. Appendix.pdf contains our prompts and human evaluation details.
Appendix
This section constitutes technical appendix (i.e. supplementary material) for the submission titled “Distilling Event Sequences from Large Language Models”.
7 Event Sequence Prompts
We now describe the specific prompt-types used (including pseudo-code for their use) to generate event sequences given individual Wikidata event concepts. We start a Wikidata event concept label and feed it into the following prompt with 3 ICL exemplars. The prompt also includes a label space of other Wikidata concepts from which we i nstruct the model to choose a concept. In case of an out-of-domain generated output ($\approx 15\%$ of total outputs), we discard those outputs.
1def build_prompt1(vocab, target_label):
2 prompt = "Use the following vocabulary to respond to the questions: " + \
[…]
6
[…]
9
[…]
12
[…]
14 f"Answer:"
15
16 return prompt
Listing 1: Initial Trigger Prompt
Following the output from the above prompt trigger, we feed the in-domain outputs to the following ICL iterative prompt with conjunctive event exemplars (i.e. X and Y). We then successively use this prompt by feeding model generated outputs as new event concepts. Restricting the generated vocabulary to existing Wikidata concepts allows for a reasonable recall evaluation even from a sparse reference set. We illustrate this approach in Figure 2 in the main paper.
1def build_prompt2(vocab, target_labels):
2
3 prompt = "Use the following vocabulary to respond to the questions: " + \
[…]
7
[…]
10
[…]
13
[…]
16
[…]
19
[…]
21 f"Answer:"
22
23 return prompt
Listing 2: Iterative ICL Prompt with Conjunctive Examples
Figure 3: Annotation interface for classifying model generated response vs reference (anonymized) for the ATOMIC event descriptions.
8 Evaluator Models (Example Outputs and Prompts)
For a model generated event sequence $\alpha=a_{1},a_{2},..,a_{n}$, we consider all possible pairs of events $(a_{i},a_{i+1})$ as (trigger, consequence) pairs where $a_{0}$ is the initial trigger sequence and $i\in{1,..,n}$ are all events generated by the model that follow $a_{0}$. Then we use the following prompt to evaluate correctness of all such possible pairs and use this as proxy for a true precision evaluation of the generated sequences.
1def build_precision_eval(trigger, consequence):
2 prompt = "Respond to the questions below with a (YES/NO) with a historical example:" + \
[…]
16 return prompt
Listing 3: Prompt used to evaluate correctness of a given event pair
As mentioned in the limitations section, a key shortcoming of our approach here is that we do not evaluate the correctness of the post-hoc explanations generated by the model for it’s classification label. We leave that analysis for future work.
9 ATOMIC Human Evaluations
Amazon Mechanical Turk (AMT) is a platform for non-expert works to perform microtasks, in our case – human annotations. Three authors with graduate degrees in computer science carried out these human evaluations. Figure 3 shows the interface provided to these human annotators. As described in the main paper, human evaluators were asked to answer the following questions–
  • Are both responses functionally similar?
  • Which response do you prefer?
  • Which response, if any, is completely irrelevant?
Functionally Similar Responses
For responses to be functionally similar, they must convey the same meaning even if there is major lexical mismatch between the actual tokens.
Human Preferences
The preferences elicited here are based on a humans degree of reasonableness given an event trigger. We observed a high Fleiss $\kappa=0.81$ indiciating a high degree of agreement between the annotators.
Irrelevant Responses
Here, again the annotators were asked to exercise judgment in what they may find a completely unreasonable response to a given event trigger.
We release the results from these human annotations for a comprehensive analysis and any additional findings that readers may infer.